OpenStudy (anonymous):

Colin is buying a used car. Colin borrows $5000 for 4 years at 12% annual interest. How much interest will Colin pay? A. $2000 B. $2200 C. $3000 D,$3200 E.None of the choices are correct

OpenStudy (mikurout):

simple interest or compound interest ?

OpenStudy (anonymous):

I don't know

OpenStudy (mikurout):

If it is simple interest, then put the formula I =P*T*R/100 I= total interest P= principle amount ($ 5000) T= Time period (4) R= rate of interest (12)

OpenStudy (anonymous):

how do you solve it can you explain me please

OpenStudy (mikurout):

R= 12% means u have to pay extra 12% of the total money for every year u keep it with u. So find out what is 12% of the total money and u have to pay that extra amount of the money at the end of each year.

OpenStudy (anonymous):

12%/100=5000/4 ???

OpenStudy (mikurout):

12% of 5000 is (12*5000)/100=600

OpenStudy (anonymous):

so the answer is E right

OpenStudy (mikurout):

this 600 is the interest of the 1st year. if u r keeping money for 4 years, then u have to give 4 times of 600 as interest.

OpenStudy (anonymous):

2400

OpenStudy (mikurout):

Yes, u got it.

OpenStudy (anonymous):

so what i have to do

OpenStudy (mikurout):

What u have to do? with what?

OpenStudy (anonymous):

so the answer is E

OpenStudy (mikurout):

yes it is E.

OpenStudy (anonymous):

thank you

OpenStudy (mikurout):

welcome..